the length of a rectangle is 7 meters more than twice the width the area is 60 square meters find the length and width
Answer by amarjeeth123(569)

Let the width be x.
Then the length is (2x+7).
Area=length*width
x(2x+7)=60
2x^2+7x-60=0
2x^2-8x+15x-60=0
2x(x-4)+15(x-4)=0
(2x+15)(x-4)=0
x=4 is the only feasible solution as x cannot be negative.
The width is 4 metres.
The length is 15 metres.